The sum of two numbers is 7 and their difference is 3. What are the two numbers?
Answer by stanbon(75887)

The sum of two numbers is 7 and their difference is 3. What are the two numbers?
----
Equations:
x + y = 7
x - y = 3
-----
Add and solve for "x":
2x = 10
x = 5
-----
Solve for "y":
x + y = 7
5 + y = 7
y = 2
----------
Cheers,
Stan H.
===========